# Frequency Decomposition-Driven Unsupervised Domain Adaptation for Remote Sensing Image Semantic Segmentation

### v1

@@ 4.6 Loss Functions
The proposed FD-GLGAN capitalizes on three loss functions, namely the segmentation loss $L_{\rm{seg}}$, the typical adversarial loss $L_{\rm{adv}}$, the frequency alignment loss $L_{\rm{f}}$, and the cross-entropy loss $L_{\rm{d}}$. These loss functions are defined as follows.
4.6.1 Generator
@@
$L_{\rm{seg}}$, $L_{\rm{f}}$ and $L_{\rm{adv}}$ are proposed to optimize the generator. In particular, $L_{\rm{seg}}$ is computed directly by images from the source domain and their corresponding labels. Mathematically, $L_{\rm{seg}}$ takes the following form:
$L_{\rm{seg}}({P}_{s})=-\lambda^{seg}\sum_{H,W}\sum_{c\in{\mathcal{C}}}Y_{s}^{(H,W,c)}\left[\mathrm{log}({P}_{s}^{(H,W,c)})\right],$ (5)
@@
where ${\mathcal{C}}$ is the category set whose cardinality is $C$ while $Y_{s}$ and $\lambda^{seg}$ denote the labels from the source domain and the weighting coefficient for the segmentation loss, respectively.
$L_{\rm{f}}$ is computed to align decomposed cross-domain features generated by multiscale HLFDs as follows:
@@
$\begin{split}&L_{\rm{f}}\left(LF_{s}^{\emph{1-4}},LF_{t}^{\emph{1-4}},HF_{s}^{\emph{1-4}},HF_{t}^{\emph{1-4}}\right)\\ &=L_{\rm{l}}\left(LF_{s}^{\emph{1-4}},LF_{t}^{\emph{1-4}}\right)+L_{\rm{h}}\left(HF_{s}^{\emph{1-4}},HF_{t}^{\emph{1-4}}\right)\\ &=\lambda^{f}\cdot\displaystyle\sum_{i=1}^{\emph{4}}\displaystyle\sum_{d=1}^{D_{i}}\frac{\left(\mathcal{P}(LF_{s}^{i,d})-\mathcal{P}(LF_{t}^{i,d})\right)}{D_{i}}+\lambda^{f}\cdot\displaystyle\sum_{i=1}^{\emph{4}}\displaystyle\sum_{d=1}^{D_{i}}\frac{\left(\mathcal{P}(HF_{s}^{i,d})-\mathcal{P}(HF_{t}^{i,d})\right)}{D_{i}},\end{split}$ (6)
@@
where $\lambda^{f}$ is a weighting coefficient, $D_{i}$ is the number of channels for the $i$-th scale, and $\mathcal{P}$ stands for the average pool, which gathers spatial information for each channel $d$. The frequency alignment loss function is designed to efficiently align the feature maps across the source and target domains.
@@
$L_{\rm{adv}}\left({I}_{t}\right)=-\lambda^{adv}\sum_{H,W}\left[\mathrm{log}\left(\mathcal{D}({I}_{t})^{(H,W,0)}\right)\right],$ (7)
where $\lambda^{adv}$ is the weighting coefficient for the adversarial loss while $\mathcal{D}$ denotes the GLDis. Note that the artificial labels 00 and $1$ are assigned to samples drawn from the source domain and the target, respectively. In Eq. (7), the label 00 in the superscript ${}^{(H,W,0)}$ is set to intentionally mislead GLDis. Therefore, the discriminator is motivated to bridge the prediction gap between the source and target domains.
@@ 4.6.2 GLDis
$L_{\rm{d}}({I}_{s},{I}_{t})=-\sum_{H,W}\left[\mathrm{log}(\mathcal{D}({I}_{s})^{(H,W,0)})+\mathrm{log}(\mathcal{D}({I}_{t})^{(H,W,1)})\right].$ (9)
4.6.3 Overall Loss Function
@@
Finally, the overall loss function proposed for FD-GLGAN is given as follows:
$\mathop{\max}_{\mathcal{D}}\ \mathop{\min}\limits_{\mathcal{G}}L(X_{s},Y_{s},X_{t}),$ (10)
@@
where $\mathcal{D}$ and $\mathcal{G}$ denote GLDis and the generator, respectively. Eq. (10) is designed to minimize the segmentation loss in the generator for the source images while maximizing the output space similarity of the source and target domains. Notably, the generator and the discriminator are trained in a classical GAN-based manner, whereas the generator and GLDis are trained alternatively. More specifically, the generator is first trained using Eq. (8) with the parameters of GLDis being held constant before GLDis is updated with Eq. (9).
5 Experiments And Discussions
5.1 Datasets Description
@@
To verify the effectiveness and robustness of the proposed FD-GLGAN, extensive experiments were conducted on two open-source benchmark datasets (Gerke, 2014) involving fine-resolution images obtained from two cities (Potsdam [^1] https://www.isprs.org/education/benchmarks/UrbanSemLab/2d-sem-label-potsdam.aspx and Vaihingen [^2] https://www.isprs.org/education/benchmarks/UrbanSemLab/2d-sem-label-vaihingen.aspx). The images in both datasets were obtained from aerial photography but of different Ground Sampling Distance (GSD). More specifically, the images of Potsdam have clearer ground objects with a GSD of $5$ cm as compared to a GSD of $9$ cm for Vaihingen. Furthermore, images in Potsdam possess richer spectrum information from four frequency bands, namely, InfraRed, Red, Green and Blue (IRRGB), whereas images in Vaihingen only three bands, namely, InfraRed, Red and Green channels (IRRG). In addition, the difference in geographical locations results in more prominent discrepancies even for the same ground object categories due to different road layouts, building types and vegetation forms. These discrepancies present a major challenge to UDA methods.
Figure 8: Image samples of size $1024\times 1024$ from (a) P-RGB, (b) P-IRRG and (c) V-IRRG, respectively. Despite some visual similarities between these two datasets, they exhibit large discrepancies in the distribution of land objects.
@@
1) Potsdam: There are $24$ very fine-resolution true orthophotos of size $6000\times 6000$ in Potsdam dataset. In the experiments, we divided these orthophotos into a training set of $18$ patches and a test set of $6$ patches as follows:
  • The training set (18 patches): {6_10, 7_10, 2_12, 3_11, 2_10, 7_8, 5_10, 3_12, 5_12, 7_11, 7_9, 6_9, 7_7, 4_12, 6_8, 6_12, 6_7, 4_11};
  • The test set (6 patches): {2_11, 3_10, 4_10, 5_11, 6_11, 7_12}.
@@
2) Vaihingen: The Vaihingen dataset contains $16$ very fine-resolution true orthophotos of size $2500\times 2000$ pixels. In our experiments, these orthophotos are divided into a training set of $12$ patches and a test set of $4$ patches as follows:
  • The training set (12 patches): {1, 3, 23, 26, 7, 11, 13, 28, 17, 32, 34, 37};
  • The test set (4 patches): {5, 21, 15, 30}.
@@
All images in both datasets are provided with semantic labels, including six ground object classes, namely, Building (Bui.), Tree (Tre.), Low Vegetation (Low.), Car, Impervious Surface (Imp.) and Clutter/Background. Fig. 5 illustrates two groups of images sampled from the Potsdam and Vaihingen datasets. Visual inspection of Fig. 5 confirms large discrepancies across the two datasets. In the sequel, we divided the Potsdam dataset into two subsets, namely, “P-IRRG” and “P-RGB”, taking into account the data size and image channels as compared to those of Vaihingen. The Vaihingen dataset forms only one dataset denoted as “V-IRRG”. By treating these three subsets as the source and target domains, we designed three UDA tasks with increasing levels of difficulty as listed in Table 2. It is worth noting that these three tasks exhibit different levels of domain shifts. For instance, Task $1$, i.e. P-IRRG to V-IRRG, suffers from the domain shift caused by large discrepancies in the spatial resolution, the ground object style and the illumination environment, though the images from both domains possess the same spectrum information. In contrast, the domain shift problem in Task $2$, i.e. P-RGB to V-IRRG, is caused by the frequency band difference in addition to all discrepancies in Task $1$. Finally, Task $3$, i.e. V-IRRG to P-RGB, encounters the same domain shift problem as Task $2$ with a greater challenge as the size of the Vaihingen dataset in the source domain is much smaller than that of the Potsdam dataset in the target domain. In summary, these three UDA tasks were designed to explore the effectiveness of the UDA method under consideration in cross-location, in-image mode, and cross-image mode UDA tasks.
Table 2: Unsupervised Domain Adaptation Tasks
@@
Table 3: Nine methods were evaluated in our following experiments, including two non-UDA and seven UDA methods.
Index Method Reference
Description
1 Source-only (Jha et al., 2019)
Images and labels from the source domain are utilized to train the segmentation network before it is tested directly on the target domain.
2 DANN (Ganin et al., 2016)
The first UDA method that applies adversarial learning to solve the domain shift problem on the image classification task.
3 AdasegNet (Tsai et al., 2018)
This UDA method performs multi-level adversarial learning with two discriminators while assuming that the source and target domains share a similar output space.
4 Advent (Vu et al., 2019)
This UDA method aligns the entropy maps of the output derived by the generator for the source and target domains.
5 GANAI (Benjdira et al., 2019)
The first UDA method designed for semantic segmentation of remote sensing images based on the GAN framework.
6 TriADA (Yan et al., 2019)
This UDA method employs a triplet branch, considering both domains simultaneously to learn a domain-invariant classifier using a domain similarity discriminator.
7 CCAGAN (Chen et al., 2022)
This UDA method develops a category-certainty attention module focusing on the misaligned regions in the category level.
8 MBATA-GAN (Ma et al., 2023b)
This UDA method develops a mutually boosted attention-based transformer to augment the high-level features and a feature discriminator to learn the transferable features.
9 Supervised (Jha et al., 2019)
Images and labels from the target domain are used to train and test the segmentation network, which stands for the optimal performance that any UDA method can expect.
5.2 Evaluation Metrics

### v2

@@ III-F Loss Functions
The proposed De-GLGAN capitalizes on four loss functions, namely the segmentation loss $L_{\rm{seg}}$, the typical adversarial loss $L_{\rm{adv}}$, the frequency alignment loss $L_{\rm{f}}$, and the cross-entropy loss $L_{\rm{d}}$. These loss functions are defined as follows.
III-F1 Generator
@@
$L_{\rm{seg}}$, $L_{\rm{f}}$ and $L_{\rm{adv}}$ are proposed to optimize the generator. In particular, $L_{\rm{seg}}$ is computed directly using images from the source domain and their corresponding labels. Mathematically, $L_{\rm{seg}}$ takes the following form:
$L_{\rm{seg}}(\bm{P}_{s})=-\lambda^{seg}\sum_{h,w}\sum_{c\in{\mathcal{C}}}\bm{Y}_{s}^{(h,w,c)}\left[\mathrm{log}(\bm{P}_{s}^{(h,w,c)})\right],$ (7)
@@
where ${\mathcal{C}}$ is the category set whose cardinality is the number of categories $C$, $\bm{Y}_{s}$ of size $\mathbb{R}^{H\times W\times C}$ denotes the labels from the source domain, and $\lambda^{seg}$ is a scalar, representing the weighting coefficient for the segmentation loss. $L_{\rm{f}}$ is computed to align decomposed cross-domain features generated by multiscale HLFDs as follows:
@@
$\begin{split}&L_{\rm{f}}\left(\bm{LF}_{s}^{\emph{1-4}},\bm{LF}_{t}^{\emph{1-4}},\bm{HF}_{s}^{\emph{1-4}},\bm{HF}_{t}^{\emph{1-4}}\right)\\ &=L_{\rm{l}}\left(\bm{LF}_{s}^{\emph{1-4}},\bm{LF}_{t}^{\emph{1-4}}\right)+L_{\rm{h}}\left(\bm{HF}_{s}^{\emph{1-4}},\bm{HF}_{t}^{\emph{1-4}}\right)\\ &=\lambda^{f}\cdot\displaystyle\sum_{i=1}^{\emph{4}}\displaystyle\sum_{d=1}^{d_{i}}\frac{\left(\mathbf{P}(\bm{LF}_{s}^{i,d})-\mathbf{P}(\bm{LF}_{t}^{i,d})\right)}{d_{i}}\\ &+\lambda^{f}\cdot\displaystyle\sum_{i=1}^{\emph{4}}\displaystyle\sum_{d=1}^{d_{i}}\frac{\left(\mathbf{P}(\bm{HF}_{s}^{i,d})-\mathbf{P}(\bm{HF}_{t}^{i,d})\right)}{d_{i}},\end{split}$ (8)
@@
where the sizes of $\bm{LF}_{s}^{i}$, $\bm{LF}_{t}^{i}$, $\bm{HF}_{s}^{i}$ and $\bm{HF}_{t}^{i}$ are $d_{i}\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}$, $d_{i}$ is the encoding dimension with $i\in\left\{1,2,3,4\right\}$ being the stage index, $\lambda^{f}$ is a scalar, representing the weighting coefficient, and $\mathbf{P}$ stands for the average pooling, which gathers spatial information for each channel $d$. The frequency alignment loss function is designed to efficiently align the feature maps across the source and target domains.
@@
$L_{\rm{adv}}\left(\bm{I}_{t}\right)=-\lambda^{adv}\sum_{h,w}\left[\mathrm{log}\left(\mathbf{D}(\bm{I}_{t})^{(h,w,0)}\right)\right],$ (9)
where $\lambda^{adv}$ is a scalar, representing the weighting coefficient for the adversarial loss while $\mathbf{D}$ denotes the GLDis. Note that the artificial labels 00 and $1$ are assigned to samples drawn from the source domain and the target domain, respectively. In Eq. (9), the label 00 in the superscript (h,w,0) is set to intentionally mislead GLDis. Therefore, the discriminator is motivated to bridge the prediction gap between the source and target domains.
@@ III-F2 GLDis
$L_{\rm{d}}(\bm{I}_{s},\bm{I}_{t})=-\sum_{h,w}\left[\mathrm{log}(\mathbf{D}(\bm{I}_{s})^{(h,w,0)})+\mathrm{log}(\mathbf{D}(\bm{I}_{t})^{(h,w,1)})\right].$ (11)
III-F3 Overall Loss Function
@@
Finally, the overall loss function $L$ proposed for De-GLGAN is given as follows:
$\mathop{\max}_{\mathbf{D}}\ \mathop{\min}\limits_{\mathbf{G}}L(\bm{X}_{s},\bm{Y}_{s},\bm{X}_{t}),$ (12)
@@
where $\bm{X}_{s}$ and $\bm{X}_{t}$ are with the size of $\mathbb{R}^{{H\times W\times Z}}$, and $\mathbf{D}$ and $\mathbf{G}$ denote GLDis and the generator, respectively. Eq. (12) is designed to minimize the segmentation loss in the generator for the source images while maximizing the output space similarity of the source and target domains. Notably, the generator and the discriminator are trained in a classical GAN-based manner, whereas the generator and GLDis are trained alternatively. More specifically, the generator is first trained using Eq. (10) with the parameters of GLDis being held constant before GLDis is updated with Eq. (11). The overall optimization process of De-GLGAN is presented in Algorithm 1 for better understanding. Finally, the well-trained generator $\hat{\mathbf{G}}$ is used for testing in the target domain.
Algorithm 1 Optimization Process of De-GLGAN
0:  Labeled source dataset $D_{s}$; Unlabeled target dataset $D_{t}$; Initialized generator $\mathbf{G}$, multiscale HLFDs and discriminator $\mathbf{D}$; Weighting coefficients $\lambda^{seg}$, $\lambda^{f}$ and $\lambda^{adv}$.
0:  Well-trained generator $\hat{\mathbf{G}}$.
1:  while in total epochs do
2:     Sample a mini-batch $\bm{X}_{s}$, $\bm{Y}_{s}$ and $\bm{X}_{t}$;
3:     Input $\bm{X}_{s}$ into $\mathbf{G}$ to obtain $\bm{F}_{s}^{\emph{1-4}}$ and $\bm{P}_{s}$;
4:     Input $\bm{X}_{t}$ into $\mathbf{G}$ to obtain $\bm{F}_{t}^{\emph{1-4}}$ and $\bm{P}_{t}$;
5:     Input $\bm{F}_{s}^{\emph{1-4}}$ and $\bm{F}_{t}^{\emph{1-4}}$ into multiscale HLFDs to obatin $*\bm{LF}_{s}^{\emph{*}}/\bm{LF}_{t}^{\emph{*}}$ and $*\bm{HF}_{s}^{\emph{*}}/\bm{HF}_{t}^{\emph{*}}$;
6:     Calculate $\bm{I}_{s}$ and $\bm{I}_{t}$ from $\bm{P}_{s}$ and $\bm{P}_{t}$ based on Eq. (5) and Eq. (6);
7:     Calculate $L_{\rm{seg}}$ based on Eq. (7);
8:     Calculate $L_{\rm{f}}$ based on Eq. (8);
9:     Fix $\mathbf{D}$ and input $\bm{I}_{t}$ into $\mathbf{D}$ to calculate $L_{\rm{adv}}$ based on Eq. (9);
10:     Update $\mathbf{G}$ based on Eq. (10);
11:     Unfix $\mathbf{D}$ and input $\bm{I}_{s}$ and $\bm{I}_{t}$ into $\mathbf{D}$ to calculate $L_{\rm{d}}$ based on Eq. (11);
12:     Update $\mathbf{D}$ based on Eq. (11).
13:  end while
14:  return  $\hat{\mathbf{G}}$ = $\mathbf{G}$.
IV Experiments And Discussions
IV-A Datasets Description
IV-A1 ISPRS Potsdam and Vaihingen
@@
This dataset [63] involves fine-resolution images obtained from two cities. The images in both datasets were obtained from aerial photography but of different Ground Sampling Distance (GSD). More specifically, the images of Potsdam have clearer ground objects with a GSD of $5$ cm as compared to a GSD of $9$ cm for Vaihingen. Furthermore, images in Potsdam possess richer spectrum information from four frequency bands, namely, InfraRed, Red, Green and Blue (IRRGB), whereas images in Vaihingen only three bands, namely, InfraRed, Red and Green channels (IRRG). All images in ISPRS Potsdam and Vaihingen are provided with semantic labels, including six ground object classes, namely, Building (Bui.), Tree (Tre.), Low Vegetation (Low.), Car, Impervious Surface (Imp.) and Clutter/Background.
@@
Potsdam: There are $24$ very fine-resolution true orthophotos of size $6000\times 6000$ in Potsdam dataset. In the experiments, we divided these orthophotos into a training set of $18$ images and a test set of $6$ images.
@@
Vaihingen: The Vaihingen dataset contains $16$ very fine-resolution true orthophotos of size $2500\times 2000$ pixels. In our experiments, these orthophotos are divided into a training set of $12$ images and a test set of $4$ images.
Fig. 8(a-c) illustrates two groups of images sampled from the Potsdam and Vaihingen datasets. Visual inspection of Fig. 8(a-c) confirms large discrepancies across the two datasets. However, since the scene categories of Potsdam and Vaihingen are similar, the category distributions are close as shown in Fig. 8(d). Therefore, the domain shift in UDA tasks constructed on Potsdam and Vaihingen mainly comes from geographical location and image mode. In the sequel, we divide the Potsdam dataset into two subsets, namely, “P-IRRG” and “P-RGB”, taking into account the data size and image channels as compared to those of Vaihingen. P-IRRG contains infrared, red, green bands while P-RGB contains red, green and blue bands. The Vaihingen dataset forms only one dataset denoted as “V-IRRG” including infrared, red, green bands. By treating these three subsets as the source and target domains, we design the first three UDA tasks with increasing levels of difficulty as listed in Table II.
Figure 8: Image samples of size $1024\times 1024$ from (a) P-RGB, (b) P-IRRG and (c) V-IRRG, respectively. Despite some visual similarities between these two datasets, they exhibit large discrepancies in the distribution of land objects. (d) Class distribution statistics for the Potsdam and Vaihingen.
@@
It is worth noting that these three tasks exhibit different levels of domain shifts. For instance, Task $1$, i.e., P-IRRG to V-IRRG, suffers from the domain shift caused by large discrepancies in the spatial resolution, the ground object style and the illumination environment, though the images from both domains possess the same spectrum information. In contrast, the domain shift problem in Task $2$, i.e., P-RGB to V-IRRG, is caused by the frequency band difference in addition to all discrepancies in Task $1$. Finally, Task $3$, i.e., V-IRRG to P-RGB, encounters the same domain shift problem as Task $2$ with a greater challenge as the size of the Vaihingen dataset in the source domain is much smaller than that of the Potsdam dataset in the target domain. In summary, these three UDA tasks were designed to explore the effectiveness of the UDA method under consideration in cross-location, in-image mode, and cross-image mode UDA tasks.
Figure 9: Image samples of size $1024\times 1024$ from (a) LoveDA Rural and (b) LoveDA Urban, respectively. (c) Class distribution statistics for the Rural and Urban scenes. The main difference between these two scenes are the large intra-class variance and the inconsistent class distributions.
IV-A2 LoveDA Rural and Urban
The LoveDA dataset contains two scenes, namely Rural and Urban. The Rural and Urban comprise $2358$ and $1833$ high-resolution optical remote sensing images, respectively. All images are with the size of $1024\times 1024$ pixels. The images provide three channels, namely Red, Green, and Blue (RGB), with a ground sampling distance of $30$ cm. The dataset encompasses seven landcover categories, including Background (Bac.), Building (Bui.), Road, Water (Wat.), Barren (Bar.), Forest (For.), and Agriculture (Agr.) [64]. These images were collected from three cities in China (Nanjing, Changzhou, and Wuhan).
Rural: The $2358$ images are divided into two parts, with $1366$ images for training and $992$ images for testing. Specifically, the training set contains images indexed from 00 to $1365$, while the test set spans images from $2522$ to $3513$.
Urban: We divided the $1833$ images into $1156$ images for training and $677$ images for testing. The former contains images indexed from $1366$ to $2521$, while the latter spans images from $3514$ to $4190$.
Fig. 9(a) and (b) present two groups of images sampled from the LoveDA Rural and Urban scenes. The discrepancy between the two scenes primarily arises from variations in ground object categories caused by geographical location. To be specific, there is intra-class variance across scene ground objects, such as shape, layout, and scale. Additionally, the class distributions are inconsistent between rural and urban scenes as shown in Fig. 9(c). These differences bring great challenges to the learning of transferable features. Following the common practice reported in the literature [64, 33], we conducted the last two UDA tasks as listed in Table II. During the training stage, the size of the sliding window was set to $256\times 256$ to collect the training samples dynamically. Therefore, we can obtain $1280$ Vaihaigen images, $13824$ Potsdam images, $37728$ Rural images and $29328$ Urban images.
TABLE II: UDA Tasks
@@ IV-A2 LoveDA Rural and Urban
Task 4 Rural Urban
Task 5 Urban Rural
TABLE III: Eleven methods were evaluated in our following experiments, including two non-UDA and nine UDA methods.
Index Method Reference Description
1 Source-only [65] Images and labels from the source domain are utilized to train the segmentation network before it is tested directly on the target domain.
2 DANN [66] The first UDA method that applies adversarial learning to solve the domain shift problem on the image classification task.
3 AdasegNet [11] This UDA method performs multi-level adversarial learning with two discriminators while assuming that the source and target domains share a similar output space.
4 Advent [12] This UDA method aligns the entropy maps of the output derived by the generator for both domains.
5 GANAI [49] The first UDA method designed for semantic segmentation of remote sensing images based on the GAN.
6 TriADA [44] This UDA method employs a triplet branch for remote sensing images, considering both domains simultaneously to learn a domain-invariant classifier using a domain similarity discriminator.
7 CCAGAN [32] This UDA method, designed for remote sensing images, develops a category-certainty attention module focusing on the misaligned regions in the category level.
8 MBATA-GAN [21] This UDA method develops a mutually boosted attention-based transformer to augment the high-level features and a feature discriminator to learn the transferable features of remote sensing images.
9 MASNet [52] This UDA method stores domain-invariant prototypical representations by employing a feature memory module, bridging the domain distribution discrepancy of two remote sensing domains.
10 DSSFNet [60] This UDA method decouples ground object features to learn domain-invariant semantic representations and domain-specific style information for remote sensing images.
11 Supervised [65] Images and labels from the target domain are used to train and test the segmentation network, which stands for the optimal performance that any UDA method can achieve.
All models in the experiments were implemented with the PyTorch framework on a single NVIDIA GeForce RTX 3090 GPU with $24$-GB RAM. Furthermore, the generator was optimized with the Stochastic Gradient Descent (SGD) optimizer and the Nesterov acceleration with a momentum of $0.9$ and a weight decay of $5\times 10^{-4}$. The initial learning rate was set to $2.5\times 10^{-4}$, decreasing with a polynomial decay of power $0.9$. For GLDis, the Adam optimizer was employed with the learning rate of $10^{-4}$ with a polynomial decay of power $0.9$. The momentum of Adam is set to $\left[0.9,0.99\right]$. In addition, the multi-level adversarial learning scheme was also adopted with parameters set to their default values derived from AdasegNet [11] in which $\lambda^{seg}$ and $\lambda^{adv}$ were set to $1.0$ and $0.001$, respectively. Finally, $\lambda^{f}$ was set to $0.01$ according to the parameter analysis that will be presented in Sec. IV-G.
IV-B Evaluation Metrics